A student ran the following reaction in the laboratory at 324 K: 2NO(g) + Br2(g) 2NOBr(g) When she introduced 0.137 moles of NO(g) and 0.119 moles of Br2(g) into a 1.00 liter container, she found the equilibrium concentration of Br2(g) to be 6.92×10-2 M. Calculate the equilibrium constant, Kc, she obtained for this reaction.
Answer #1

2NO + Br2 ----------> 2NOBr

0.137 M 0.119M 0 ( at t = 0 )

0.137-x 0.119-x 2x

But [Br2] = 6.92*10^-2 = 0.0692

i.e. 0.119 - x = 0.0692, So x = 0.0498

Thus, [NO] = 0.137 - x = 0.137 - 0.0498 = 0.0872 M

[NOBr] = 2x = 2*0.0498 = 0.0996 M

Kc = [NOBr]^2 / ([NO]^2 * [Br2]

Kc = (0.0996)^2 / ((0.0872)^2 * 0.0692)

Kc = 0.00992 / ( 0.007604 * 0.0692 )

Kc = 18.85 ....... Answer